kalynne55:

Eric works for an airline, and he needs to calculate the weight of passengers and luggage before takeoff. The number of passengers is in cell A1. The airline uses an average weight of 180 pounds per person, so Eric enters 180 in cell A2. The weight of the checked luggage is entered in cell A3. Each passenger is alloted 10 pounds for carry-on luggage, so he enters 10 in cell A4. Which formula should Eric use?

Laylalyssa:

Welcome to QC! can you post the answer choices please?

Laylalyssa:

@kalynne55

kalynne55:

yes one sec

kalynne55:

=A1*A2*A3*A4 =A1*A2+A3+A1*A4 =A1*(A2+A3+A1)*A4 =A1*A2+(A3+A1)*A4

Laylalyssa:

Do you have any ideas on what it might be?

kalynne55:

the 3rd one

Laylalyssa:

how

Laylalyssa:

first u have to multiply A1 nd A2, to get the weight of the passengers.

kalynne55:

ok

supie:

What you have to do is \[ A1*A2=weight\ of\ passengers\] \[A1*A4= Weight\ of\ luggage\] Now we need to add the checked weight. That is an A3. So now we add A3 with those So we would get \[A1*A2 + A1*A4 + A3\]

kalynne55:

thnx yall that helped so much

Laylalyssa:

ur welcome

supie:

np!